OpenStudy (ashleythho):

Given: 5x + 3 > 4x + 7. Choose the graph of the solution set.

OpenStudy (3mar):

Need help? @ashleythho

Angle:

316cce6b25334ca591c27a4b94b3004e.png http://prntscr.com/drjm65 I am assuming that this is the graph that you are referring to :)

Angle:

In any case, when given an inequality that looks like 5x + 3 > 4x + 7 the first task would always be to "solve" the equation that is, to get x by itself on one side

Angle:

if the > is confusing you, you can think of it as a = when you are solving this inequality I will do it this way so 5x + 3 = 4x + 7 let's first get all the terms that have an x to one side soooo moving 4x from the right side, to the left side means that we should subtract 4x from both sides 5x + 3 = 4x + 7 -4x -4x


5x - 4x + 3 = 4x - 4x + 7 then combine like terms would get you x + 3 = 7

Angle:

from here, in order to get x alone, we want to move 3 to the other side to do that we would subtract 3 from both sides x + 3 = 7 becomes -3 -3 - - - - - - - - x = 4

Angle:

now, since at the beginning I had replaced the > with a = we can put it back in now that we've solved it :) so we get x > 4

Angle:

the > means that (stuff) "Is greater than" (other stuff) greater than means that it is on the right side on the number line (because the bigger numbers are there so x > 4 means that it would be shaded to the right of 4